# Fourfold path to Thermality: Inequivalent purifications of Rindler wedge

### v1

@@ IV.2 Particle content along path 2
Hence, the left-moving modes in wedge $\mathrm{R}_{2}$ remain in vacuum when mapped to a wedge $\mathrm{R}_{3}$; no particle creation occurs in this region. This is consistent with the preservation of the positive-frequency structure of the modes under the coordinate transformation.
We now examine the particle content in region $\mathrm{R}_{3}$ as seen from the vacuum defined in region $\mathrm{R}_{1}$, by following the composite path $R_{1}\to R_{2}\to R_{3}$ (Path 3). The state in $\mathrm{R}_{1}$ is taken to be the Rindler vacuum, satisfying
$\overrightarrow{\hat{a}}_{1}|0_{R_{1}}\rangle=0\qquad and\qquad\overleftarrow{\hat{b}}_{1}|0_{R_{1}}\rangle=0,$ (132)
Along this path, the left-moving modes experience Bogoliubov mixing from $\mathrm{R}_{1}$ to $\mathrm{R}_{2}$, while the right-moving modes undergo mixing from $\mathrm{R}_{2}$ to $\mathrm{R}_{3}$.
Composing Bogoliubov Coefficients:
To analyze the full transition from $\mathrm{R}_{1}$ to $\mathrm{R}_{3}$, we compose the Bogoliubov coefficients using the substitution of Eq. (57) into Eq. (56), yielding:
$\begin{split}f_{j}(\Omega)=\int d\omega\int d\Omega\Big[\alpha_{ji}(\Omega,\omega)\big(\alpha_{ji}^{*}(\Omega,\omega)\,f_{j}(\Omega)-\beta_{ji}(\Omega,\omega)\\ \,f_{j}^{*}(\Omega)\big)+\beta_{ji}(\Omega,\omega)\big(\alpha_{ji}(\Omega,\omega)\,f_{j}^{*}(\Omega)-\beta_{ji}^{*}(\Omega,\omega)\,f_{j}(\Omega)\big)\Big],\end{split}$
Rearranging terms and comparing with the standard expansion of modes in terms of annihilation and creation operators, we identify the composed Bogoliubov coefficient:
$\beta_{31}(\nu,\omega)=\int d\Omega\left[\beta_{32}(\nu,\Omega)\,\alpha_{21}(\Omega,\omega)-\alpha_{32}(\nu,\Omega)\,\beta_{21}(\Omega,\omega)\right].$ (133)
Right-Moving mode:
We now examine the right-moving sector. In the second leg of the path, $R_{2}\to R_{3}$, the Bogoliubov transformation for right-moving modes is nontrivial, while the transformation from $R_{1}\to R_{2}$ contributes no additional right-moving excitations. Therefore, using Eqs. (100), (101), and (133) for $\Omega,\omega>0$, the integration simplifies, and we obtain:
$\overrightarrow{\beta_{31}}(\nu,\omega)=\overrightarrow{\beta_{32}}(\nu,\omega),$ (134)
Since $\overrightarrow{\beta_{32}}$ corresponds to thermal particle production (as shown previously), it follows that $\overrightarrow{\beta_{31}}$ also exhibits thermal behavior.
Left-Moving mode:
Conversely, for left-moving modes, the thermal contribution arises during the transition from $\mathrm{R}_{1}$ to $\mathrm{R}_{2}$, while the transformation from $\mathrm{R}_{2}$ to $\mathrm{R}_{3}$ preserves left-moving vacuum structure. As a result, from Eq. (133) in the left-moving case, we find:
$\overleftarrow{\beta_{31}}(\nu,\omega)=-\overleftarrow{\beta_{21}}(\nu,\omega),$ (135)
Again, since $\overleftarrow{\beta_{21}}$ is thermal, this confirms that the left-moving modes in region $\mathrm{R}_{3}$ also appear excited with thermal weight when viewed from the vacuum in $R_{1}$.
We therefore conclude that, along path 3 from region $\mathrm{R}_{1}$ to $\mathrm{R}_{3}$, both the left- and right-moving modes undergo Bogoliubov transformations that excite particles relative to the initial vacuum $\mid 0_{R_{1}}\rangle$.The left-moving mode acquires thermal excitation through the transition $R_{1}\to R_{2}$, while the right-moving sector becomes thermal due to the transformation $R_{2}\to R_{3}$, as demonstrated by Eqs. (134) and  (135).
The resulting state in wedge $\mathrm{R}_{3}$, as expressed in the number basis of left- and right-moving modes, thus takes the form:
$\mid\psi_{R_{3}}\rangle=\left(\sum_{n}^{\infty}\sqrt{p_{n}}\;\mid\overleftarrow{n}\rangle\right)\otimes\left(\sum_{m}^{\infty}\sqrt{q_{m}}\;\mid\overrightarrow{m}\rangle\right),$ (136)
Equation (136) represents a product of two pure states in the Fock space of region $\mathrm{R}_{3}$, each expressed as a thermal-like superposition of number eigenstates in the left- and right-moving modes. Although the coefficients $p_{n}$ and $q_{m}$ follow thermal occupation distributions, the resulting state is a pure vector in the single-wedge Hilbert space and does not correspond to a mixed Gibbs thermal density matrix. In the present construction, both left- and right-moving modes are supported entirely within a single Rindler wedge and generate the same local field algebra. Since no causally complementary region is traced over, there is no bipartite decomposition of the Hilbert space and hence no entanglement between Rindler wedges along the null shift. Consequently, the selectively excited state remains pure, despite exhibiting a thermal particle spectrum(Bogoliubov-induced superposition of number eigenstates) in one chiral sector.
Importantly, this state is not a purification of a thermal density matrix, nor is it structurally equivalent to a thermofield double, which involves entanglement between left and right modes. Instead, the structure of Eq. (136) arises from independent Bogoliubov excitations in the left- and right-moving modes along the composite path $R_{1}\to R_{3}$. As a result, the state is a pure [2] and direct tensor product, on the single-wedge Fock space of $\mathrm{R_{3}}$. A detailed discussion of the structure and interpretation of this state—including its relation to thermal density matrices and purification—is provided in Appendix D.
IV.4 Particle content along the horizon $U_{1}$ and $V_{2}$ axes
In this section, we analyze the Bogoliubov transformations and particle content along Path 4, wherein the Rindler wedge $\mathrm{R}_{4}$ is obtained from $\mathrm{R}_{1}$ via a null shift along the $U_{1}$-axis, followed by the transition $\mathrm{R}_{4}\to\mathrm{R}_{3}$ through a null shift along the $V_{2}$-axis, as depicted in Fig. 2.
@@ IV.4 Particle content along the horizon $U_{1}$ and $V_{2}$ axes
Since the method for computing Bogoliubov coefficients has already been established in the previous section, we directly write down the expressions for the transition $\mathrm{R}_{1}\to\mathrm{R}_{4}$, which follows the same computational steps. The Bogoliubov coefficients for the right-moving modes are given by:
@@
This confirms that no particle creation occurs in the left-moving sector during the transition from $R_{1}\to R_{4}$.
We conclude that the null shift along the $U_{1}$-axis selectively excites only the right-moving modes in $\mathrm{R}_{4}$, while the left-moving modes remain in the vacuum.
@@
$\overleftarrow{\beta^{*}_{34}}(\nu,\rho)=-\sqrt{\frac{\nu}{\rho}}\int_{-\infty}^{\infty}\frac{dv_{3}}{2\pi}\;e^{i\nu v_{3}}\;e^{i\rho v_{4}},$ (149)
@@
$\overleftarrow{\beta^{*}_{34}}(\nu,\rho)=-\frac{1}{2\pi a}\sqrt{\frac{\nu}{\rho}}\frac{2^{\frac{i(\rho+\nu)}{a}}\;\Gamma[\frac{i\nu}{a}]\;\Gamma[\frac{-i(\rho+\nu)}{a}]}{\Gamma[\frac{-i\rho}{a}]},$ (151)
@@
${\mid\overleftarrow{\beta_{34}}(\nu,\rho)\mid}^{2}=\overleftarrow{\beta^{*}_{34}}(\nu,\rho)\;\overleftarrow{\beta_{34}}(\nu,\rho),$ (152)
@@
${\mid\overleftarrow{\beta_{34}}(\nu,\rho)\mid}^{2}=\frac{1}{4\pi^{2}a^{2}}\;\frac{\nu}{\rho}\;\frac{{\mid\Gamma[\frac{i\nu}{a}]\mid}^{2}\;{\mid\Gamma[\frac{i(\nu+\rho)}{a}]\mid}^{2}}{{\mid\Gamma[\frac{i\rho}{a}]\mid}^{2}},$ (153)
@@
${\mid\overleftarrow{\beta_{34}}(\nu,\rho})\mid^{2}=\frac{1}{4\pi a}\frac{sinh(\frac{\pi\rho}{a})}{(\nu+\rho)\;sinh(\frac{\pi\nu}{a})\;sinh[\frac{\pi(\nu+\rho)}{a}]},$ (154)
Since $\mathrm{R}_{4}$ is obtained through a null shift from $\mathrm{R}_{1}$, the resulting quantum field configuration exhibits selective thermalisation, and the right-moving modes become thermally populated. In contrast, the left-moving modes remain in the vacuum, as established earlier. The corresponding quantum state in the region $\mathrm{R}_{4}$ is a tensor product state in the Hilbert space $\mathcal{H}_{R_{4}}$, and is given by[31, 32].
$\mid\psi_{R_{4}}\rangle=\left(\sum_{n=0}^{\infty}\sqrt{p_{n}}\mid\overrightarrow{n}\rangle\right)\otimes\mid\overleftarrow{0}\rangle,$ (155)
[…]
We now compute the expectation value of the number operator for left-moving modes in the region $\mathrm{R}_{3}$, to the quantum state in region $\mathrm{R}_{4}$. That is:
$\langle\overleftarrow{\hat{N}}_{34}(\nu)\rangle=\langle\psi_{R_{4}}\mid\overleftarrow{\hat{b}_{3}^{\dagger}}(\nu)\;\overleftarrow{\hat{b}}_{3}(\nu)\mid\psi_{R_{4}}\rangle,$ (156)
Substituting the Bogoliubov transformation relations from Eqs. (63) and  (64) into Eq. (156) and using the (j=3 and i=4 and also $\Omega=\nu$ and $\omega=\rho$ ),we obtain:
$\begin{split}\langle\overleftarrow{\hat{N}}_{34}(\nu)\rangle=\langle\psi_{R_{4}}\mid\int_{0}^{\infty}d\rho\int_{0}^{\infty}d\rho^{\prime}\;\overleftarrow{\beta^{*}_{34}}(\nu,\rho)\;\overleftarrow{\beta_{34}}(\nu,\rho^{\prime})\\ \overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\psi_{R_{4}}\rangle,\end{split}$ (157)
To evaluate this expression, recall from Eq. (155) that the state $\mid\psi_{R_{4}}\rangle$ is a tensor product with the left-moving mode in the vacuum. Therefore, the two-point function reduces to a vacuum expectation value:
$\langle\psi_{R_{4}}\mid\;\overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\psi_{R_{4}}\rangle={\langle\overleftarrow{0}}\mid\overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\overleftarrow{0}\rangle,$ (158)
Since the left-moving modes are in vacuum in $\mid\psi_{R_{4}}\rangle$, we obtain:
${\langle\overleftarrow{0}}\mid\overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\overleftarrow{0}\rangle=\delta(\rho-\rho^{\prime}),$ (159)
@@
This corresponds to a Planckian spectrum at a temperature of $T=\frac{a}{2\pi}$. Thus, we conclude that the left-moving modes connecting regions $\mathrm{R}_{4}$ and $\mathrm{R}_{3}$ are thermally populated with a Planckian distribution, as perceived by observers in $\mathrm{R}_{3}$, when the field is selectively excited $|\psi_{R_{4}}\rangle$.
We now define the state vector in the Rindler region $\mathrm{R}_{3}$ as observed from the vacuum state in $\mathrm{R}_{1}$, following the composite path $R_{1}\to R_{4}\to R_{3}$ (Path 4). The vacuum in $\mathrm{R}_{1}$ is defined by:
$\overrightarrow{\hat{a}}_{1}\mid 0_{R_{1}}\rangle=0\qquad and\qquad\overleftarrow{\hat{b}}_{1}|0_{R_{1}}\rangle=0,$ (163)
Along this path, the right-moving modes become thermally populated in the transition $R_{1}\to R_{4}$, and the left-moving modes undergo thermal excitations during the transition $R_{4}\to R_{3}$.
Composing Bogoliubov Coefficients:
As in Section C, we compose the Bogoliubov coefficients for the two-step transformation $R_{1}\to R_{4}\to R_{3}$.
$\beta_{31}(\nu,\omega)=\int d\rho\left[\beta_{34}(\nu,\rho)\,\alpha_{41}(\rho,\omega)-\alpha_{34}(\nu,\rho)\,\beta_{41}(\rho,\omega)\right].$ (164)
This relates the Bogoliubov coefficient between modes in $\mathrm{R}_{3}$ and $\mathrm{R}_{1}$ to the intermediate transformations through region $\mathrm{R}_{4}$, following the same structure as Eq. (133) in the previous section.
Right-Moving mode:
In the first leg of the path, $R_{1}\to R_{4}$, the right-moving modes are thermally excited, while the second leg, $R_{4}\to R_{3}$, leaves the right-moving modes unchanged. Using Eqs. (142),  (143), and Eq (164), we obtain:
$\overrightarrow{\beta_{31}}(\nu,\omega)=-\overrightarrow{\beta_{41}}(\nu,\omega),$ (165)
Since, $\overrightarrow{\beta_{41}}$ is thermal, it follows that $\overrightarrow{\beta_{31}}$ also exhibits thermal behavior.
Left-moving mode:
For the left-moving mode, the thermal contribution arises from the transformation $R_{4}\to R_{3}$, while the transition $R_{1}\to R_{4}$ leaves the left-moving modes in the vacuum. Therefore,
$\overleftarrow{\beta_{31}}(\nu,\omega)=\overleftarrow{\beta_{34}}(\nu,\omega),$ (166)
Since $\overleftarrow{\beta_{34}}$ is thermal, this confirms that the left-moving modes in region $\mathrm{R}_{3}$ also appear excited with thermal weight when viewed from the vacuum in $\mathrm{R}_{1}$.
We therefore conclude that, along the path 4 from region $\mathrm{R}_{1}$ to $\mathrm{R}_{3}$, both the left- and right-moving modes undergo Bogoliubov transformations that excite particles relative to the initial vacuum $\mid 0_{R_{1}}\rangle$.The left-moving mode acquires thermal excitation through the transition $R_{1}\to R_{4}$, while the right-moving sector becomes thermal due to the transformation $R_{4}\to R_{3}$, as demonstrated by Eqs. (165) and  (166).
The resulting state in wedge $\mathrm{R}_{3}$, as expressed in the number basis of left- and right-moving modes, thus takes the form:
$\mid\psi_{R_{3}}\rangle=\left(\sum_{m}\sqrt{q_{m}}\mid\overleftarrow{m}\rangle\right)\otimes\left(\sum_{n}\sqrt{p_{n}}\mid\overrightarrow{n}\rangle\right),$ (167)
Eq. (167) describes a product of two pure states in the Fock space of region $\mathrm{R}_{3}$ as in the previous section, each written as a thermal-like superposition over number eigenstates. Although the coefficients $p_{n}$ and $q_{m}$ exhibit thermal distributions, the global state is pure and unentangled between the left- and right-moving modes.
This structure is not equivalent to a thermofield double, which would involve entanglement across the two modes. Rather, the thermal character in each mode arises independently due to the composite Bogoliubov transformation along the path $R_{1}\to R_{4}\to R_{3}$. A detailed discussion of the structure and interpretation of this state is provided in Appendix D.
IV.5 Bogoliubov Transformation for two observes in Rindler spacetime
Figure 3: Rindler spacetime with an observer experiencing acceleration a and g, shown in green and red, respectively.
$T=\frac{e^{a\;x_{1}}}{a}\;\sinh\;(a\;t_{1}),$ (168)
$X=\frac{e^{a\;x_{1}}}{a}\;\cosh(a\;t_{1}).$ (169)
which in light-cone coordinates takes the form:
$U=T-X=-\frac{e^{-a\;u_{1}}}{a},$ (170)
$V=T+X=\frac{e^{a\;v_{1}}}{a}.$ (171)
Similarly, for Rindler coordinates with acceleration $g$, we have:
$T=\frac{e^{g\;x_{2}}}{g}\;\sinh\;(g\;t_{2}),$ (172)
$X=\frac{e^{g\;x_{2}}}{g}\;\cosh(g\;t_{2}),$ (173)
and in light-cone form:
$U=T-X=-\frac{e^{-g\;u_{2}}}{g},$ (174)
$V=T+X=\frac{e^{g\;v_{2}}}{g},$ (175)
$u_{1}=\frac{-1}{a}\;ln\bigg(\frac{a}{g}\bigg)+\frac{g}{a}\;u_{2},$ (176)
$v_{1}=\frac{1}{a}\;ln\bigg(\frac{a}{g}\bigg)+\frac{g}{a}\;v_{2},$ (177)
Substituting Eq. (176) into Eq. (178), we obtain:
By the integration technique, we obtained the following result.
Similarly, the $\beta^{*}$-coefficient becomes
Hence, by integrating Eq. (182), we get
$\overrightarrow{\beta^{*}}_{\Omega,\omega}=0,$ (183)
Following a similar analysis for left-moving modes, we obtain:
$\overleftarrow{\beta^{*}}_{\Omega,\omega}=0,$ (185)
V Conformal Transformations and Energy Flux in Two-Dimensional Conformal Field Theories
@@
V.1 Two-Dimensional CFT on Rindler Spacetime in the Minkowski Vacuum
To investigate the Minkowski Vacuum $|0_{M}\rangle$, as perceived by a uniformly accelerated Rindler observer via different paths. We are using the Schwarzian derivative to compute the expectation value of the components of the energy-momentum tensor through the Minkowski Vacuum $|0_{M}\rangle$. In a 2d CFT on Minkowski spacetime, the energy-momentum tensor satisfies[5]:
$\partial_{\mu}T^{\mu\nu}=0,$ (186)
From the above equation, we have the non-vanishing components are $T_{U_{M}U_{M}}$ and $T_{V_{M}V_{M}}$. Furthermore, in the Minkowski Vacuum state $|0_{M}\rangle$,
$\langle 0_{M}|T_{U_{M}U_{M}}|0_{M}\rangle=0\qquad\langle 0_{M}|T_{V_{M}V_{M}}|0_{M}\rangle=0,$ (187)

### v2

@@ IV.2 Particle content along path 2
Hence, the left-moving modes in wedge $\mathrm{R}_{2}$ remain in vacuum when mapped to a wedge $\mathrm{R}_{3}$; no particle creation occurs in this region. This is consistent with the preservation of the positive-frequency structure of the modes under the coordinate transformation. So both the left-moving and right-moving modes have a thermal distribution of particles
IV.4 Path 4: Particle content along the shifts along $U_{1}$ and $V_{2}$ axes
@@
In this section, we analyze the Bogoliubov transformations and particle content along Path 4, wherein the Rindler wedge $\mathrm{R}_{4}$ is obtained from $\mathrm{R}_{1}$ via a null shift along the $U_{1}$-axis, followed by the transition $\mathrm{R}_{4}\to\mathrm{R}_{3}$ through a null shift along the $V_{2}$-axis, as depicted in Fig. 2. Since the method for computing Bogoliubov coefficients has already been established in the previous section, we directly write down the expressions for the transition $\mathrm{R}_{1}\to\mathrm{R}_{4}$, which follows the same computational steps. The Bogoliubov coefficients for the right-moving modes are given by:
@@
This confirms that no particle creation occurs in the left-moving sector during the transition from $R_{1}\to R_{4}$. We conclude that the null shift along the $U_{1}$-axis selectively excites only the right-moving modes in $\mathrm{R}_{4}$, while the left-moving modes remain in the vacuum. We therefore have,
$|\overleftarrow{0_{k}\rangle}_{R_{1}}=|\overleftarrow{0_{k}\rangle}_{R_{4}},$ (139)
@@
$\overleftarrow{\beta^{*}_{34}}(\nu,\rho)=-\sqrt{\frac{\nu}{\rho}}\int_{-\infty}^{\infty}\frac{dv_{3}}{2\pi}\;e^{i\nu v_{3}}\;e^{i\rho v_{4}}.$ (145)
@@
$\overleftarrow{\beta^{*}_{34}}(\nu,\rho)=-\frac{1}{2\pi a}\sqrt{\frac{\nu}{\rho}}\frac{2^{\frac{i(\rho+\nu)}{a}}\;\Gamma[\frac{i\nu}{a}]\;\Gamma[\frac{-i(\rho+\nu)}{a}]}{\Gamma[\frac{-i\rho}{a}]}.$ (147)
@@
${\mid\overleftarrow{\beta_{34}}(\nu,\rho)\mid}^{2}=\overleftarrow{\beta^{*}_{34}}(\nu,\rho)\;\overleftarrow{\beta_{34}}(\nu,\rho).$ (148)
@@
${\mid\overleftarrow{\beta_{34}}(\nu,\rho)\mid}^{2}=\frac{1}{4\pi^{2}a^{2}}\;\frac{\nu}{\rho}\;\frac{{\mid\Gamma[\frac{i\nu}{a}]\mid}^{2}\;{\mid\Gamma[\frac{i(\nu+\rho)}{a}]\mid}^{2}}{{\mid\Gamma[\frac{i\rho}{a}]\mid}^{2}}.$ (149)
@@
${\mid\overleftarrow{\beta_{34}}(\nu,\rho})\mid^{2}=\frac{1}{4\pi a}\frac{sinh(\frac{\pi\rho}{a})}{(\nu+\rho)\;sinh(\frac{\pi\nu}{a})\;sinh[\frac{\pi(\nu+\rho)}{a}]}.$ (150)
Since $\mathrm{R}_{4}$ is obtained through a null shift from $\mathrm{R}_{1}$, the resulting quantum field configuration exhibits selective thermalisation, and the right-moving modes become thermally populated. In contrast, the left-moving modes remain in the vacuum, as established earlier.
We now compute the expectation value of the number operator for left-moving modes in the region $\mathrm{R}_{3}$, to the quantum state in region $\mathrm{R}_{4}$. Substituting the Bogoliubov transformation relations from Eqs. (64) and  (65) into Eq. (148) and using the (j=3 and i=4 and also $\Omega=\nu$ and $\omega=\rho$ ),we obtain:
$\begin{split}\langle\overleftarrow{\hat{N}}_{34}(\nu)\rangle=\langle\overrightarrow{0}_{R_{4}}\mid\int_{0}^{\infty}d\rho\int_{0}^{\infty}d\rho^{\prime}\;\overleftarrow{\beta^{*}_{34}}(\nu,\rho)\;\overleftarrow{\beta_{34}}(\nu,\rho^{\prime})\\ \overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\overrightarrow{0}_{R_{4}}\rangle,\end{split}$ (151)
The vacuum expectation value becomes:
$\langle\overrightarrow{0}_{R_{4}}\mid\;\overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\overrightarrow{0}_{R_{4}}\rangle={\langle\overleftarrow{0}}\mid\overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\overleftarrow{0}\rangle,$ (152)
Since the left-moving modes are in vacuum, we obtain:
${\langle\overleftarrow{0}}\mid\overleftarrow{\hat{b}}_{4}(\rho^{\prime})\;\overleftarrow{\hat{b}}_{4}^{\dagger}(\rho)\mid\overleftarrow{0}\rangle=\delta(\rho-\rho^{\prime}).$ (153)
@@
This corresponds to a Planckian spectrum at a temperature of $T=\frac{a}{2\pi}$. Thus, we conclude that the left-moving modes connecting regions $\mathrm{R}_{4}$ and $\mathrm{R}_{3}$ are thermally populated with a Planckian distribution, as perceived by observers in $\mathrm{R}_{3}$.
V Particle content via Virasoro anomaly
@@
V.1 Analysis of the four paths via Virasoro Anomaly
V.2 Path 1
We obtain the particle content in $R_{3}$ via path 1. To investigate the Minkowski Vacuum $|0_{M}\rangle$, as perceived by a uniformly accelerated Rindler observer via different paths. We are using the Schwarzian derivative to compute the expectation value of the components of the energy-momentum tensor through the Minkowski Vacuum $|0_{M}\rangle$. We have the non-vanishing components are $T_{U_{M}U_{M}}$ and $T_{V_{M}V_{M}}$. Furthermore, in the Minkowski Vacuum state $|0_{M}\rangle$,
$\langle 0_{M}|T_{U_{M}U_{M}}|0_{M}\rangle=0,\qquad\langle 0_{M}|T_{V_{M}V_{M}}|0_{M}\rangle=0.$ (157)